OpenStudy (anonymous):

Find the square. (2x – 6)2 A. 4x2 – 24x + 36 B. 4x2 – 8x + 36 C. 4x2 + 36 D. 4x2 – 12x + 36

OpenStudy (anonymous):

Lets first expand the brackets to see what we get. (2x - 6)^2 = (2x - 6)(2x - 6) Ok now lets find the square using the FOIL method which is what you may have learnt (2x - 6)(2x - 6) 2x * 2x = 4x^2(First terms) 2x * -6 = -12x(Outside terms) -6 * 2x = -12x(Inside terms) -6 * -6 = 36(Last terms) Now lets write out what we got 4x^2 - 12x - 12x + 36 And then we simplify by collecting like terms 4x^2 - 24x + 36 Thats out final answer, so your answer is A.

OpenStudy (anonymous):

(2x-6) times (2x-6) 2x(2x)=4x sq. 2x(-6)=-12x -6(2x)=-12x -6(-6)=36 answer 4x^2-24x+36 -12+-12=-24